Can someone please help me with this problem?
Determine if the following equation is symmetric with respect to the y-axis, x-axis, the origin, or none of these.
a. y=3x^2
b. x^2+4y^2=16

Answer by lwsshak3(11628)

Determine if the following equation is symmetric with respect to the y-axis, x-axis, the origin, or none of these.
a. y=3x^2
This is an equation of a parabola that takes the form: x^2=4py, which is a parabola that opens upwards. In such a parabola we get the same y-coordinate at x or -x which makes the graph symmetric with respect to the y axis
..
b. x^2+4y^2=16
x^2/16+y^2/4=1
This is an equation of an ellipse with center at the origin (0,0) This makes the graph symmetric with respect to both the y and x axes